### v2

@@ 4.2.2 Bayesian models of sentence processing
A second advantage of the internal neural dynamics of DFT is that they generate novel predictions, e.g., by relating the interaction of representations to their metric distance in feature space. In general, interaction is excitatory for more similar representations, and inhibitory for more dissimilar representations. These are the dynamics which allow the stabilization of activation peaks; neural recordings consistent with these dynamics have been observed in cat visual cortex, for example (Jancke et al., 1999). With respect to language, it has been observed that different interpretations of homophonous words—which are more distal in semantic feature space—tend to inhibit each other, leading to slowed response times, while different interpretations of polysemous words—which are more proximal in semantic feature space—tend to prime each other, leading to faster response times (Frisson, 2015; Klepousniotou and Baum, 2007; Klepousniotou et al., 2012; MacGregor et al., 2015; Rodd et al., 2002). From the current perspective, these effects can be seen as a natural consequence of the basic dynamics of lateral interaction in neural fields, analogous to effects of metric feature distance on target-distractor interaction observed in the domains of eye saccades (Kopecz and Schöner, 1995), manual reaching movements (Erlhagen and Schöner, 2002), and speech articulation (Tilsen, 2009). Careful variation of the feature distance between interpretations in priming studies could potentially be used to empirically constrain the parameters of lateral interaction which determine the radii of excitatory and inhibitory projection in neural fields governing processing on semantic dimensions.
@@ 4.3 Conclusion
Using a single theoretical framework (DFT) and mathematics of description (differential equations) allows explicit integration across the cognitive and sensory-motor domains. Previous DFT modeling work has linked neural representations of conceptual structure (Jackendoff, 2002) with visual perception (Grieben and Schöner, 2022) and visual search behavior (Sabinasz et al., 2023). Explicit coupling between abstract cognitive processes and sensation/movement has been termed “grounding” of cognition (e.g., Sabinasz and Schöner, 2023; Sabinasz et al., 2023). As described in Section 1.5.1, existing DFT models of speech and language have focused on phonetic dimensions of articulatory movement and auditory perception, or perceptual dimensions of objects like color, size, and spatial position. In this paper we have extended DFT to linguistic semantics through the lens of lexical polysemy, a cognitive domain that appears relatively distinct from sensation and behavior. Our use of the same theoretical framework and mathematics as these previous models paves the way for explicitly coupling the cognitive and sensory-motor aspects of language. This would represent a significant step towards a complete neurocognitive model of language linking meaning and form: a grounded model of linguistic cognition.
@@ 6 Funding sources
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Appendix A Experimental stimuli

### v3

@@ 4.2.2 Bayesian models of sentence processing
A second advantage of the internal neural dynamics of DFT is that they generate novel predictions, e.g., by relating the interaction of representations to their metric distance in feature space. In general, interaction is excitatory for more similar representations, and inhibitory for more dissimilar representations. These are the dynamics which allow the stabilization of activation peaks; neural recordings consistent with these dynamics have been observed in cat visual cortex, for example (Jancke et al., 1999). With respect to language, it has been observed that different readings of homophonous words—which are more distal in semantic space—tend to inhibit each other, leading to slowed response times, while different readings of polysemous words—which are more proximal in semantic space—tend to prime each other, leading to faster response times (Frisson, 2015; Klepousniotou and Baum, 2007; Klepousniotou et al., 2012; MacGregor et al., 2015; Rodd et al., 2002). From the current perspective, these effects can be seen as a natural consequence of the basic dynamics of lateral interaction in neural fields, analogous to effects of metric feature distance on target-distractor interaction observed in the domains of eye saccades (Kopecz and Schöner, 1995), manual reaching movements (Erlhagen and Schöner, 2002), and speech articulation (Tilsen, 2009). Careful variation of the semantic distance between readings in priming studies could potentially be used to empirically constrain the parameters of lateral interaction which determine the radii of excitatory and inhibitory projection in neural fields governing processing on semantic dimensions.
@@ 4.3 Conclusion
Using a single theoretical framework (DFT) and mathematics of description (differential equations) allows explicit integration across the cognitive and sensory-motor domains. Previous DFT modeling work has linked neural representations of conceptual structure (Jackendoff, 2002) with visual perception (Grieben and Schöner, 2022) and visual search behavior (Sabinasz et al., 2023). Explicit coupling between abstract cognitive processes and sensation/movement has been termed “grounding” of cognition (e.g., Sabinasz and Schöner, 2023; Sabinasz et al., 2023). As described in Section 1.5.1, existing DFT models of speech and language have focused on phonetic dimensions of articulatory movement and auditory perception, or perceptual dimensions of objects like color, size, and spatial position. In this paper we have applied DFT to model linguistic semantics through the lens of lexical polysemy, a cognitive domain that appears relatively distinct from sensation and behavior. Our use of the same theoretical framework and mathematics as these previous models paves the way for explicitly coupling the cognitive and sensory-motor aspects of language. This would represent a significant step towards a more integrated neurocognitive model of language linking meaning and form: a grounded model of linguistic cognition.
@@ 6 Funding sources
This work was supported by a research award from Yale University to the Neurolinguistics Lab.
Appendix A Model equations and descriptions
The dynamics of the have node are given in Eq. 2.
The rate of change of activation is negatively related to current activation , defining a dynamical system with a point attractor at . represents external input to the node, e.g. from perception or intention, and represents normally distributed noise weighted by a parameter . When there is external input to the node, the node’s activation is attracted to . When there is no input, activation is attracted to 0 . is a time constant, with higher values corresponding to slower rates of evolution. In our simulations, is set to , and is set to . We set the magnitude of external input depending on the condition being simulated.
The dynamics of each of the two DNFs are given in Eq. 3.
Activation is defined for each neuron representing the relevant semantic dimension at each moment in time . Activation in the DNF has a point attractor at . The resting level is assumed to be below zero for all neurons, by convention at . We set the time constant for the DNFs to .
Each input , , and is represented as a separate Gaussian distribution of the form
where controls the amplitude of the input, controls the position of the input in the field, and controls the width of the input distribution. In our simulations, we set the amplitude of external input depending on the condition being simulated. represents input from the have node, whose amplitude is defined straightforwardly as a linear function of the activation of the node: . This is a simplification relative to most DFT models, where the amplitude of input from a node to a field would be a more complex (sigmoidal) function of node activation. Our motivation for eliminating the sigmoidal gating function on node-to-field input comes from work on lexical neighborhood effects on articulation, where non-selected lexical items (nodes) exert some influence on phonetic planning (DNFs) (Stern and Shaw, 2023b). This issue is largely orthogonal to the present study because there is only one node in the model. We chose to use the same node-field coupling dynamics as in Stern and Shaw (2023b) for the sake of simplicity and consistency. The same results could likely be achieved with a relatively shallow or “soft” sigmoidal gating function.
The amplitude of each input from one DNF to the other is given by
Appendix B Experimental stimuli